# KAN-ODEs: Kolmogorov-Arnold Network Ordinary Differential Equations for Learning Dynamical Systems and Hidden Physics

### v1

@@ III.2 Modeling Hidden Physics: Fisher-KPP PDE
In this section, we introduce the capability of KAN-ODEs to learn the hidden physics in PDEs. Key features demonstrated here are the flexibility of KAN-ODEs to be paired with higher-complexity solvers on arbitrary temporal grids, and the capability of KAN-ODEs to extract hidden symbolic functional relationships from training data. We demonstrate this using the Fisher-KPP equation representing a convection-diffusion system,
@@
$u(x,0)$ $=\frac{1}{2}\left[\tanh\left(\frac{x-0.4}{0.02}\right)-\tanh\left(\frac{x-0.6}{0.02}\right)\right].$ (14)
Note that this problem setup was adopted from [3].The spatial grid was was discretized with $\Delta x=0.04$, and the equation was solved with a time step of $\Delta t=0.5$. The obtained solution field $u(x,t)$ is shown in Fig. 4(A) and used as training data.
Figure 4: Fisher-KPP equation: (A) Solution field $u(x,t)$ of ground truth. (B) Solution field $u(x,t)$ of prediction with a learned KAN-ODE model. (C) Loss function. (D) Learned hidden physics of the reaction source term in the Fisher-KPP equation and its symbolic form.
To model the unknown reaction source term in Eq. 12, we formulate the KAN-ODE such that
@@
To utilize the backpropagation with the adjoint sensitivity method, Eq. 15 was discretized using a central difference scheme and converted to an ODE system using the Method of Lines. The KAN here was constructed with a single layer comprising a single node, [1,1,5].
The prediction with a learned model is illustrated in Fig. 4(B), showing accurate reconstruction of the solution field $u(x,t)$ after as few as 5,000 updates (Fig. 4(C)) and demonstrating that a single learned activation function comprised of five gridded basis functions ([1,1,5]) is sufficient to represent the reaction source term. This resultant activation function is shown in Fig. 4(D). We further propose expressing this single activation function symbolically, as was introduced in [21] as a feature of KAN layers. Symbolic regression on this single KAN-ODE activation using SymbolicRegression.jl [29] given the candidate symbolic expressions of $[+,-,*,/,\sin,\cos,\exp]$ returns the following expression,
@@
Given its single-layer, single-node construction, this symbolic expression is the KAN-ODE’s approximation of the true reaction source term in Eq. 12. Knowing $r$=1.0, we see that the KAN-ODE’s derived expression is remarkably close to the original formulation. Thus the KAN-ODE was not only capable of learning the dynamics of physics hidden within a PDE submodel from measurable quantities, but also of returning an accurate and human-interpretable symbolic function that mimics the true governing law with excellent accuracy.
III.3 Data-Driven Solutions: Burgers’ Equation
Our third example explores the use of a KAN-ODE to infer the unknown hidden states $u\left(x,t\right)$ of a PDE system, effectively representing all of its spatiotemporal dynamics. Let us consider the Burgers’ reaction-diffusion equation as an example.
@@ III.3 Data-Driven Solutions: Burgers’ Equation
The computational domain is defined as $x\in[-1.0,1.0]$ and $t\in[0.0,1.0]$.
In order to prepare training data, Eq. 17 was discretized using a central difference scheme with $\Delta x=0.05$ such that $\mathbf{u}\left(t\right)=\left[u\left(x_{0},t\right),u\left(x_{1},t\right),...,u\left(x_{N},t\right)\right]$. Then, the discretized PDE was solved by an ODE integrator given the initial condition of $u(x,0)=-\sin(\pi x)$. The resultant solution field is illustrated in Fig. 5(A). The solution profiles at the selected times ($t\in\{0.1,0.3,0.5,0.7,0.9\}$) were collected as training data.
To learn the latent solution, we model the KAN-ODEs such that
@@
In this formulation, the KAN works as a non-linear operator for the partial differential equation (PDE) solution. The KAN-ODE was constructed with 2 layers: [51, 10, 5], [10, 51, 5]. Ten nodes in the intermediate layer were used to preserve the high-dimensional information. Note that the numbers of the input and output dimensions of the KAN are determined by the number of discretized state variables in $\mathbf{u}$. The KAN parameters were updated with a learning rate of 0.01 with the ADAM optimizer.
@@
Figure 5: Burgers’ equation: (A) Solution fields $u(x,t)$ for ground truth and prediction with an initial guess and the learned parameters. The black dashed lines in the solution field of ground truth indicates the training data. (B) Loss function. (C) Training data and prediction at times $t$= 0.1, 0.3, 0.5, 0.7, and 0.9 s. (D) Ground truth and inferred solution profiles at selected times $t$= 0.2, 0.4, 0.6, 0.8, and 1.0 s.
@@
After 20,000 training epochs (Fig. 5(B)), the KAN-ODE model can be seen in Fig. 5(A) to predict the temporal evolution of the PDE well despite only receiving five training snapshots. These snapshots and their accurate KAN-ODE dynamical representations are shown in Fig. 5(C), where strong agreement is seen at all spatial locations, even those near the shock wave where the training data sampling is very sparse. More importantly, the trained KAN-ODE model is able to extrapolate effectively to infer the solution fields at the unseen testing times in Fig. 5(D), once again with strong performance at the shock. As a whole, this case demonstrates the ability of the KAN-ODE to not only learn the dynamics of the Burgers’ convection-diffusion PDE from velocity data, but also to leverage its temporal grid agnosticism to accurately predict the system’s behavior at unseen times.
@@ IV Conclusions
This work proposed the use of Kolmogorov-Arnold networks as gradient-getters in the Neural ODE framework. This novel combination of data-driven techniques advances our capabilities in the field of scientific machine learning to efficiently infer interpretable and modular dynamical system models from various sources of data. The replacement of MLPs with KANs in this framework preserves the black-box style approach with zero prior knowledge of the dynamical system physics needed for training. However, it opens the door to significant interpretability of the learned model (”opening” the black box) via activation function visualization and symbolic regression. Quantitatively, the KAN-ODE framework is shown to outperform similar MLP-based Neural ODEs when given the same training task. The strong neural scaling offered by the Kolmogorov-Arnold representation holds when applied as a KAN-ODE, allowing for significantly better training and testing performance in a fraction of the computational time. We additionally demonstrated the capability of the KAN-ODE to extrapolate information from sparse or limited temporal data to the rest of a solution domain, and to learn symbolic expressions for PDE source terms with high accuracy via optional postprocessing. In addition to these improvements over MLP-based Neural ODEs, KAN-ODEs offers an alternative to SINDy, PINN, and similar interpretable machine learning techniques in its flexibility and generality - no prior knowledge or assumptions are needed about the governing laws or candidate function space to train the KAN-ODE system effectively, although such information can be added if desired as shown here in the source term regression of the Fisher-KPP PDE. KAN-ODEs are a compact, efficient, flexible, and interpretable approach for data-driven modeling of dynamical systems that show significant promise for use in the field of scientific machine learning.
@@ Acknowledgement
Appendix A. KAN-ODE Performance for Different KAN Architectures
Additional tests were performed to provide the KAN-ODE performance for varying KAN architectures in the examples of KPP-Fisher equation and Burgers’ equation. Table A1 shows the test matrix with different sizes of the KAN. For the example of KPP-Fisher equation, a relatively small KAN was used to model the hidden physics in the interpretable form while a larger size of the KAN was constructed to build a surrogate model that predicts the entire solution fields. Figure A1 illustrates the training losses with respect to the number of parameters. Surprisingly, in the presented cases with two demonstration examples, a small size of the KAN can successfully learn a dynamical system with a training loss smaller than 10-6. Furthermore, KAN-ODEs has a scaling curve in most cases such that $\mathcal{L}\propto N^{-4}$ where $N$ is the total number of parameters in the KAN. This result empirically supports the validity of the scaling law in the original KAN [16].
Table A1: KAN-ODEs architecture tested for performance evaluation with the Fisher KPP equation and the Burgers’ equation.
@@ Appendix A. KAN-ODE Performance for Different KAN Architectures
Figure A1: Performance test with different KAN-ODEs architectures: Training loss as a function of the total number of parameters. (A) KPP Fisher equation. (B) Burgers’ equation. Scaling of $\mathcal{L}\propto N^{-4}$, where $N$ is the total number of parameters in the KAN, is also illustrated as a reference line.

### v2

@@ III.2 Modeling Hidden Physics in PDEs: Fisher-KPP PDE
In this section, we introduce the capability of KAN-ODEs to learn the hidden physics in PDEs. Key features demonstrated here are the flexibility of KAN-ODEs to be paired with higher-complexity solvers on arbitrary temporal grids, and the capability of KAN-ODEs to directly extract hidden symbolic functional relationships from training data. We demonstrate this using the Fisher-KPP equation representing a reaction-diffusion system,
@@
$u(x,0)$ $=\frac{1}{2}\left[\tanh\left(\frac{x-0.4}{0.02}\right)-\tanh\left(\frac{x-0.6}{0.02}\right)\right],$ (15)
@@
$\frac{\partial u}{\partial x}\left(0,t\right)$ $=\frac{\partial u}{\partial x}\left(1,t\right).$
Note that this problem setup was adopted from [9]. The spatial grid was discretized with $\Delta x=0.04$, and the equation was solved with a time step of $\Delta t=0.5$. The obtained solution field $u(x,t)$ is shown in Fig. 5(A) and used as training data.
Figure 5: Fisher-KPP equation: (A) Solution field $u(x,t)$ of the ground truth. (B) Solution field $u(x,t)$ of the prediction with a learned KAN-ODE model. (C) Loss function. (D) Learned hidden physics of the reaction source term in the Fisher-KPP equation and its symbolic form.
To model the unknown reaction source term in Eq. 14, we formulated the KAN-ODE such that
@@
To utilize backpropagation with the adjoint sensitivity method, Eq. 17 was discretized using a central difference scheme and converted to an ODE system using the Method of Lines. The KAN was constructed with a single layer comprising a single node, [1,1,10].
The prediction with a learned model is illustrated in Fig. 5(B), showing accurate reconstruction of the solution field $u(x,t)$ after as few as 5,000 updates (Fig. 5(C)) and demonstrating that a single learned activation function comprised of ten gridded basis functions ([1,1,10]) is sufficient to represent the reaction source term (see Appendix B for a detailed discussion of accuracy with varying KAN-ODE architectures). This resultant activation function is shown in Fig. 5(D). We further propose expressing this single activation function symbolically, as was introduced in [26] as a feature of KAN layers. Symbolic regression on this single KAN-ODE activation using SymbolicRegression.jl [42] given the candidate symbolic expressions of [+,,×$[+,-,\times,/,\sin,\cos,\exp]$ returned the following expression,
@@
Given its single-layer, single-node construction, this symbolic expression is the KAN-ODE’s approximation of the true reaction source term in Eq. 14. Knowing $r=1.0$, we see that the KAN-ODE’s derived expression is remarkably close to the original formulation. Thus, the KAN-ODE was not only capable of learning the dynamics of physics hidden within a PDE submodel from measurable quantities but also of returning an accurate and human-interpretable symbolic function that mimics the true governing law with excellent accuracy.
III.3 Data-Driven Solutions of PDEs
III.3.1 Burgers’ Equation
Our third example explores the use of a KAN-ODE to infer the unknown hidden states $u\left(x,t\right)$ of a PDE system, effectively representing all of its spatiotemporal dynamics. Here we considered the Burgers’ convection-diffusion equation as an example.
@@ III.3.1 Burgers’ Equation
The computational domain was defined as $x\in[-1.0,1.0]$ and $t\in[0.0,1.0]$.
In order to prepare training data, Eq. 19 was discretized using a central difference scheme with $\Delta x=0.05$ such that $\mathbf{u}\left(t\right)=\left[u\left(x_{0},t\right),u\left(x_{1},t\right),...,u\left(x_{N},t\right)\right]$. Then, the discretized PDE was solved by an ODE integrator given the initial and boundary conditions such that
$u\left(x,0\right)$ $=-\sin(\pi x),$ (20)
$u\left(-1,t\right)$ $=0,$ (21)
$u\left(1,t\right)$ $=0.$
[…]
@@
In this formulation, the KAN works as a non-linear operator for the partial differential equation (PDE) solution. The KAN-ODE was constructed with 2 layers: [51, 10, 5], [10, 51, 5]. Ten nodes in the intermediate layer were used to preserve the high-dimensional information. Note that the numbers of the input and output dimensions of the KAN were determined by the number of discretized state variables in $\mathbf{u}$. The KAN parameters were updated with a learning rate of 0.01 with the ADAM optimizer.
@@
Figure 6: Burgers’ equation: (A) Solution fields $u(x,t)$ for the ground truth, prediction with the initialized model, and prediction with the trained model. The black dashed lines in the solution field of the ground truth subplot indicate the training data. (B) Loss function. (C) Training data and inferred solution profiles at times $t$= 0.1, 0.3, 0.5, 0.7, and 0.9 s. (D) Unseen ground truth data and inferred solution profiles at selected times $t$= 0.2, 0.4, 0.6, 0.8, and 1.0 s.
@@
After 20,000 training epochs (Fig. 6(B)), the KAN-ODE model can be seen in Fig. 6(A) to predict the temporal evolution of the PDE well despite only receiving five training snapshots. These snapshots and their accurate KAN-ODE dynamical representations are shown in Fig. 6(C), where strong agreement is seen at all spatial locations, even those near the shock wave where the training data spatial sampling is very sparse (in addition to sparse temporal sampling). More importantly, the trained KAN-ODE model is able to extrapolate effectively to infer the solution fields at the unseen testing times in Fig. 6(D), once again with strong performance at the shock. As a whole, this case demonstrates the ability of the KAN-ODE to leverage its temporal grid agnosticism to accurately predict a system’s behavior at unseen times, and to extrapolate significant meaning from spatially and temporally sparse datasets. An in-depth discussion of the model accuracy and KAN architecture can be found in Appendix B.
III.3.2 Schrödinger Equation
The last example demonstrates surrogate modeling with KAN-ODEs in a more involved dynamical system with complex-valued states. The Schrödinger equation is a classical field equation used in quantum mechanics.
$i\frac{\partial u}{\partial t}+\frac{1}{2}\frac{\partial^{2}u}{\partial x^{2}}+\lvert u\rvert^{2}u=0,$ (23)
where $u\in\mathbb{C}^{1}$. The time and spatial domains were set to $x\in[-5,5]$ and $t\in[0,\pi/2]$. The initial condition and boundary conditions were defined as following in Eqs. 24-25. Note that the problem settings used in this example were identical with those in [6].
$u\left(x,0\right)$ $=\text{sech}\left(x\right),$ (24)
$u\left(-5,t\right)$ $=u\left(5,t\right),$ (25)
$\frac{\partial u}{\partial x}\left(-5,t\right)$ $=\frac{\partial u}{\partial x}\left(5,t\right).$
The KAN-ODE as a surrogate model (Eq. 22) was constructed with 2 layers: [402, 10, 10], [10, 402, 10]. The input dimension was doubled from the dimension of the discretized $u$ (201 to 402) to account for both the real and imaginary values of $u$.
Figure 7: Schrödinger equation: (A) Observed data $u(x,t)=\text{Re}(u)+i\text{Im}(u)$ at the selected times of $t$ = 0.1, 0.3, 0.5, 0.7, 0.9, 1.1, 1.3, and 1.5 sec as training data. Blue circles denote $\text{Re}(u)$ and red circles denote $\text{Im}(u)$. (B) Reconstructed $u(x,t)$. (C) Ground truth $u(x,t)$. Note that $\text{Re}(u)$ and $\text{Im}(u)$ denote the real and imaginary parts of $u$, respectively. Also, $\lvert u\rvert=\sqrt{\text{Re}(u)^{2}+\text{Im}(u)^{2}}$.
We once again gave the KAN-ODE training data at just eight selected times in Fig. 7(A), and yet the trained KAN-ODE successfully reconstructs the complete spatiotemporal profile of the wave field $u(x,t)$ (Fig. 7(B)), which agrees with the ground truth $u(x,t)$ (Fig. 7(C)) at all times. We find this result remarkable, especially as the KAN-ODE had zero knowledge of the underlying physics and instead was forced to infer the complete solution to the nonlinear and complex dynamics present in this system from eight samples only.
In this section, we demonstrated the robustness of KAN-ODEs by learning shock formation behavior and complex-valued wave propagation in data-lean frameworks and without prior knowledge of the system dynamics. In both cases the KAN-ODEs were able to match the training data and generalize well in the significant windows between training data slices, showing promise as a novel data-driven dynamical system modeling approach. With these PDE examples, we showed that KAN-ODEs can retrieve not only temporal dynamics but also spatial information without special treatment in the KAN architecture.
@@ IV Conclusions
This work proposed the use of Kolmogorov-Arnold networks as gradient-getters in the Neural ODE framework. This novel combination of data-driven techniques advances our capabilities in the field of scientific machine learning to efficiently infer interpretable and modular dynamical system models from various sources of data. The replacement of MLPs with KANs in this framework preserves the black-box style approach with zero prior knowledge of the dynamical system physics needed for training. However, it opens the door to significant interpretability of the learned model (”opening” the black box) via activation function visualization and symbolic regression. Quantitatively, the KAN-ODE framework was also shown to outperform similar MLP-based Neural ODEs in a wide array of metrics including training accuracy, testing accuracy, convergence speed, neural scaling rate, and generalizability. The strong neural scaling offered by the Kolmogorov-Arnold representation held when applied as a KAN-ODE with RBF basis functions, allowing for significantly better training and testing performance in a fraction of the computational time and with a significantly smaller network size. We additionally demonstrated the capability of KAN-ODEs to extrapolate information from sparse or limited temporal data to the rest of a solution domain even in stiff and complex-valued nonlinear PDEs, and to learn symbolic expressions for PDE source terms with high accuracy via optional postprocessing. In addition to these improvements over MLP-based Neural ODEs, KAN-ODEs offer an alternative to SINDy, PINN, and similar interpretable machine learning techniques in their flexibility and generality - no prior knowledge or assumptions are needed about the governing laws or candidate function space to train the KAN-ODE system effectively, although such information can be added if desired as shown here in the source term regression of the Fisher-KPP PDE and the symbolic KAN leveraged for Lotka-Volterra dynamics. KAN-ODEs are a compact, efficient, flexible, and interpretable approach for data-driven modeling of dynamical systems that show significant promise for use in the field of scientific machine learning.
@@ Acknowledgement
CRediT authorship contribution statement
Benjamin C. Koenig: Methodology, Software, Investigation, Writing - Original Draft. Suyong Kim: Conceptualization, Methodology, Software, Investigation, Writing - Original Draft. Sili Deng: Funding Acquisition, Resources, Writing - Review &\&& Editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix A. Sparse Regression on Lotka-Volterra KAN-ODE
Symbolic functions corresponding to the Symbolic KAN-ODE plotted in Fig. 4(B) are provided in Table A1. See Fig. 4(B) for a visual representation of the twelve activations and of the hidden layer nodes $n_{1}$ through $n_{3}$. The selection of candidate functions was carried out to balance accuracy with sparsity. For example, a linear term suffices to capture the linear behavior of $l1$ connecting $x$ and $n_{1}$, while $l2$ connecting $x$ and $n_{2}$ requires an additional $1/x$ term due to its strong nonlinear behavior. Similarly, the relatively complex $l11$ connecting $n_{3}$ to $dx/dt$ uses a linear term in addition to a polynomial division term. All activations were derived independently based on the input/output pairs present in the training data.
Table A1: Symbolic KAN for Lotka-Volterra dynamics. The first layer takes the input of $x$ and $y$ while the second layer takes the input of $n_{1}$, $n_{2}$, and $n_{3}$. Note that $n_{i}(x,y)=n_{i}(x)+n_{i}(y)$ and $(dx/dt)(n_{1},n_{2},n_{3})=(dx/dt)(n_{1})+(dx/dt)(n_{2})+(dx/dt)(n_{3})$ by definition in the KAN.
Layer Activation Symbolic Expression
1 $n_{1}\left(x\right)$ $0.204\displaystyle 0.545x-$ 0.204
$n_{2}\left(x\right)$ $-0.277x+0.425+\frac{0.794}{x}$
$n_{3}\left(x\right)$ $0.635\displaystyle 0.334x-$ 0.635
$n_{1}\left(y\right)$ $-0.605y+0.220-\frac{0.120}{y}$
$n_{2}\left(y\right)$ $-0.506y+1.864-\frac{0.108}{y}$
$n_{3}\left(y\right)$ $-0.780y-0.102$
2 $\frac{dx}{dt}\left({n_{1}}\right)$ $0.411\displaystyle 0.090n_{1}^{3}+0.520n_{1}^{2}+0.890n_{1}-$ 0.411
$\frac{dy}{dt}\left({n_{1}}\right)$ $0.748\displaystyle 0.168n_{1}^{3}-0.887n_{1}^{2}+2.242n_{1}+$ 0.748
$\frac{dx}{dt}\left({n_{2}}\right)$ $-{n_{2}^{2}}+4.990n_{2}+\frac{0.439n_{2}}{n_{2}^{2}+0.0695}-1.652$
$\frac{dy}{dt}\left({n_{2}}\right)$ $-n_{2}+\frac{-3.201n_{2}}{n_{2}^{2}+0.719n_{2}+0.675}$
$\frac{dx}{dt}\left({n_{3}}\right)$ $0.490\displaystyle 0.684n_{3}+\frac{2.929n_{3}}{n_{3}^{2}+1.307n_{3}+1.41}-$ 0.490
$\frac{dy}{dt}\left({n_{3}}\right)$ $-\frac{1.281n_{3}}{n_{3}^{2}+0.800}+0.353$
Appendix B. Varying KAN-ODE Architectures for Fisher-KPP and Burgers’ Equations
Additional tests were performed to provide KAN-ODE performance metrics for varying KAN architectures in the Fisher-KPP equation and Burgers’ equation examples. Table B1 shows the test matrix with different sizes of KANs. For the Fisher-KPP equation, a relatively small KAN with very few parameters was used to model the hidden physics in the interpretable form (i.e. replacing only the source term), while larger KANs with more parameters (thanks to the larger input/output dimension) were constructed for the Burgers’ equation to build a surrogate model that predicted the entire solution field. Figure B1 illustrates the training losses with respect to the number of parameters in these two examples. We surprisingly found in both cases that small architectures of KANs, i.e. those with zero to one hidden node and three to five grid points, can successfully learn these dynamical systems with a training loss smaller than 10-6. Furthermore, in most cases, KAN-ODEs have a scaling rate of loss $\mathcal{L}\propto N^{-4}$, where $N$ is the total number of parameters in the KAN. This result empirically supports the validity of the scaling law in the original KAN [26], as well as results seen in the Lotka-Volterra example shown in Fig. 3(C).
Table B1: KAN-ODEs architecture tested for performance evaluation with the Fisher-KPP equation and the Burgers’ equation.
@@ Appendix B. Varying KAN-ODE Architectures for Fisher-KPP and Burgers’ Equations
Figure B1: Performance test with different KAN-ODE architectures: Training loss as a function of the total number of parameters. (A) Fisher-KPP equation. (B) Burgers’ equation. Loss scaling of $\mathcal{L}\propto N^{-4}$, where $N$ is the total number of parameters in the KAN, is also illustrated as a reference line.
Appendix C. Additional Example with the Allen-Cahn Equation
In this appendix, we investigated the Allen-Cahn equation which describes phase separation in multi-component alloy systems. The Allen-Cahn equation was used as another canonical example to demonstrate two KAN-ODE tasks: modeling hidden physics and reconstructing complete solution fields. The Allen-Cahn equation is expressed by Eq. C1.
$\frac{\partial u}{\partial t}=0.0001\frac{\partial^{2}u}{\partial x^{2}}+5u-5u^{3}.$ (C1)
The computational domain was defined as $x\in[-1,1]$ and $t\in[0,1]$. The initial and boundary conditions were set as Eqs. C2-C3.
$u\left(x,0\right)$ $=x^{2}\cos\left(\pi x\right),$ (C2)
$u\left(-1,t\right)$ $=u\left(1,t\right),$ (C3)
$\frac{\partial u}{\partial x}\left(-1,t\right)$ $=\frac{\partial u}{\partial x}\left(1,t\right).$
Equation C1 was discretized with $\Delta x$ = 0.05 using a central difference scheme. Then, the discretized PDE was solved by the Tsit5 ODE integrator.
C.1. Modeling Hidden Physics in PDEs
To model the unknown source term in Eq. C1, we formulated the KAN-ODE such that
$\frac{\partial u}{\partial t}=0.0001\frac{\partial^{2}u}{\partial x^{2}}+\text{KAN}\left(u,\bm{\theta}\right).$ (C4)
Similarly to the Fisher-KPP example, Eq. C4 was discretized using a central difference scheme and converted to an ODE system using the Method of Lines. We constructed the KAN model with a single layer comprising a single node, [1,1,10], thereby encoding just one activation function.
Figure C1: Allen-Cahn equation: (A) Solution field $u(x,t)$ of ground truth. (B) Solution field $u(x,t)$ of prediction with a learned KAN-ODE model. (C) Loss function. (D) Learned hidden physics of the reaction source term in the Allen-Cahn equation and its symbolic expression.
The KAN-ODE model was trained on the data shown in Fig. C1(A). In this example, we applied an early stop when the error reached 10-7. The prediction with the learned model is illustrated in Fig. C1(B), showing accurate reconstruction of the solution field $u(x,t)$ after $\sim$65,000 updates (Fig. C1(C)). This resultant single activation function is shown in Fig. C1(D). For the modeled activation function, a symbolic expression was derived, using the SymbolicRegression.jl package [42]. Given the candidate symbolic expressions of [+,,×,/][+,-$\times$,/][ + , - , × , / ], it returned the expressions tabulated in Table. C1. Among them, the following expression showed good accuracy,
Table C1: Seven candidate symbolic expressions of the reaction source term in the Allen-Cahn equation (Eq. C1). The complexities of the symbolic equations and the corresponding losses are also tabulated.
Complexity Loss Equation
1 9.1219$\times$10-1 $u$
3 6.5775$\times$10-1 $1.8526u$
5 6.5775$\times$10-1 $u/0.5398+8.3105\times 10^{-4}$
7 4.2106$\times$10-1 $(2.4821-u^{2})u$
9 1.1523$\times$10-5 $(5.0015-5.0021u^{2})u$
11 1.0832$\times$10-5 $5.0021(0.9999-u^{2})u$
13 1.0493$\times$10-5 $5.0021u(0.9999-u(u+3.7283\times 10^{-4}))$
$\text{KAN}\left(u\right)=(5.0015-5.0021u^{2})u.$ (C5)
Equation C5 is remarkably close to the exact source term of $5u-5u^{3}$. Note that all the symbolic expressions with errors on the order of 10-5 in Table C1 have an identical form with slightly different coefficients. This example provides another demonstration of accurate and human-interpretable modeling of hidden physics from measurable quantities via KAN-ODEs.
C.2. Data-Driven Solutions of PDEs
We further demonstrate KAN-ODEs as surrogate models using the Allen-Cahn equation by replacing the entire right-hand side of Eq. C1 with a KAN. To mimic sparse data measurements, only five profiles at the selected times $t\in\{0.1,0.3,0.5,0.7,0.9\}$ were collected from the ground truth simulation and used as training data (Fig. 22(A)), while testing involved reconstruction of the entire time history. The KAN-ODE as a surrogate model was constructed using 2 layers with 10 hidden nodes ([41,10,10], [10,41,10]). Equation 22 was trained to learn the dynamics of the phase transition as shown in the loss plot (Fig. C2(B)). The resulting field reconstruction of $u(x,t)$ is illustrated in Fig. C2(C), which qualitatively matches well with the ground truth. The spatio-temporal error between the KAN-ODE model and the ground truth ($u^{\text{KAN}}(x,t)-u^{\text{true}}(x,t)$) in Fig. C2(D) depicts near-zero error at and immediately near each training time slice, and small but noticeable error in the large windows between time steps, especially near the end of the simulation time. For example, the error is not appreciable at time $t$ = 0.1, but begins to grow before shrinking back to near-zero by $t$ = 0.3 when the second observation takes place. This result implies that the KAN-ODE needs more training data in order to extrapolate quantitatively well to times significantly outside of its training data window, like $t$ = 1.5 or $t$ = 2.0. It does, however, highlight the strong capability of KAN-ODEs to accurately model and reconstruct physical phenomena between training samples with very sparse measurements and without any prior physical knowledge included.
Figure C2: Allen-Cahn equation: Demonstration of KAN-ODEs as a surrogate model. (A) Ground truth $u(x,t)$. The dashed lines indicate the training data at $t$ = 0.1, 0.3, 0.5, 0.7, 0.9. (B) MSE loss during training. (C) Reconstructed $u(x,t)$ by the trained KAN-ODE model. (D) The error between the reconstructed $u^{\text{KAN}}(x,t)$ and the ground truth $u^{\text{true}}(x,t)$.